OpenStudy (anonymous):

need help! substitution method 4x + y = 16 7x + 3y = 9

OpenStudy (anonymous):

so far I have these steps completed... 4x + y =16 4x + y = 16 + 4x 7x 3(16 + 4) = 9 7x + 48 + -12) = 9

syed98 (syedmohammed98):



syed98 (syedmohammed98):

I hope u get it :)

OpenStudy (anonymous):

so I put the 5x in the first equation? @SyedMohammed98

syed98 (syedmohammed98):

nah 39 = 5x so divide both sides by 5 in order to get the value of x so u get x = 7.8

syed98 (syedmohammed98):

insert value of x (7.8) in any equation to find the value of y :)

OpenStudy (anonymous):

NOT UNDERSTANDING...Can you write it in steps for me...If you can start from where I left off at the top of the screen

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@dan815 @ParthKohli